The temperatures of the junctions of a bismuth-silver thermocouple are maintained at 0°C and 0.001°C. Find the thermo-emf (Seebeck emf) developed. For bismuth-V/deg^2.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. θ = 0.001°C a = – 46 × 10^–6 v/deg, b = –0. 48 × 10^–6 v/deg^2 Emf = a base BlAg θ +(1/2) b base BlAg θ^2 = – 46 × 10^–6 × 0.001 – (1/2) × 0.48 × 10^–6 (0.001)^2 = – 46 × 10^–9 – 0.24 × 10^–12 = – 46.00024 × 10^–9 = – 4.6 × 10^–8 V